34.62 mL of a 0.3 M Mg(OH)2 solution are needed to neutralize a 15-mL solution of HCl on unknown concentration. Calculate the concentration of HCl solution.
1
Expert's answer
2020-06-12T07:28:57-0400

Mg(OH)2+2HCl=MgCl2+2H2O

1.Trace by finding the amount of Mg(OH)2. To do this, we multiply the molarity by the volume (liters):

0.03462litres *0.3 mol/litres= 0.010386 mol Mg(OH)2

2.based on the reaction we find the amount of HCl.that is, HCl is 2 times more than Mg(OH)2, so multiplying Mg(OH)2 by 2 gives the amount of HCl: 0.010386*2=0.020472mol

3.Find the concentration of HCl by dividing the amount of substance by the volume (liters): 0.020472mol/0.015 litres=1.3648mol/litres

Answer:1.3648mol/litres